2. Suppose Swift Jet Airlines estimates that it’s no-show rate is approximately 9%. What is the probability that no passenger will be bumped if the airline booked 230 passengers on a 200-seat plane?
Answer #1
n= 230 p= 0.9100
here mean of distribution=μ=np= 209.3
and standard deviation σ=sqrt(np(1-p))= 4.3402
for normal distribution z score =(X-μ)/σx
therefore from normal approximation of binomial distribution and continuity correction:

probability that no passenger will be bumped if the airline booked 230 passengers on a 200-seat plane :

probability = P(X<200.5) = P(Z<-2.03)= 0.0212